# MAGELLAN: Metacognitive predictions of learning progress guide autotelic LLM agents in large goal spaces

### v2

@@ D.3.2 Goal sampling strategies
We analyze the sampling strategies of the different methods. In Figure 15, it appears that MAGELLAN has a sampling strategy close to the one of EK-Online-ALP. The main difference is that EK-Online-ALP does not select impossible goals thanks to the use of expert knowledge. MAGELLAN’s sampling strategy quickly begins to sample goals of type ”Grow herbivores” after sampling goals of type ”Grow plants”, which contrasts with the strategy of Online-ALP. Indeed, Online-ALP mostly selects goals already encountered, getting stuck on one goal type before moving to another. The similarity between MAGELLAN and EK-Online-ALP indicates that MAGELLAN is able to cluster the goal space dynamically, using this information to generalize its competence estimation to sample interesting goals not yet discovered. More results on the dynamic clustering of the goal space by MAGELLAN are given in Appendix D.4.2.
@@ D.4.1 Per-goal Success Probability estimation on test set
In this section, we provide a detailed analysis of the results presented in Section 4.3, specifically examining the ability to generalize success probability estimations to test goals across different goal types (see Figure 16). As expected, Online-ALP exhibits the largest errors, as it can only assign a success probability of 0 to unseen goals. In contrast, both MAGELLAN and EK-Online-ALP achieve highly accurate estimations for ”Grasp”, ”Grow plants”, and ”Grow herbivores”. However, both methods tend to overestimate the generalization abilities of the policy on ”Grow carnivores” goals.
@@ D.4.2 Evolution of the embeddings
  • After mastering the goals of type ”Grasp”, Figure 17(b) shows several clustered appeared. All ”Grasp” goals are in the same cluster with a high estimated success probability zone. The goals of type ”Grow plants” are also clustered together and close to the zone of high estimated success probability. That is a hint that MAGELLAN has already picked them as the next candidate type for the curriculum. It also correctly places the ”Grow plants” goals from the test set in the same cluster as the ones from the train set. However, it still mixes them with impossible goals, underlying that it does not fully master this type of goal. The ”Grow herbivores” and ”Grow carnivores” are mixed with other impossible goals.
  • After achieving mastery of ”Grow plants”, in Figure 17(c), we see both ”Grasp” and ”Grow plants” are correctly clustered in the zone of high success probability, with the goals from the test set correctly placed into the two clusters. The goals from the type ”Grow herbivores” and ”Grow carnivores” are clustered together and almost separated from the impossible goals. Their cluster is close to the zone of high estimated probability.
  • Once ”Grow herbivores” has been solved, we see in Figure 17(d) that the three achieved goal types are clustered in three groups, all in the zone of high estimated success probability. The ”Grow carnivores” are far from this zone but still clustered together, apart from the impossible ones. It also appears that the model places ”Grow carnivores” goals from the test close to the ”Grow herbivores”.
  • Finally, when the type ”Grow carnivores” is mastered, in Figure 17(e), all the goals from the possible goal types (from both train and test) are placed in the zone of high success probability. The goals of type ”Grow X” form three clusters tightly packed. The ”Grasp” type is separated from the other types.
@@
(c) ”Grow plants” type of goals mastered.
(d) ”Grow herbivores” type of goals mastered.
(e) ”Grow carnivores” type of goals mastered.
@@ D.4.3 Embedding of impossible goals
As detailed in Section A.3, a goal is considered impossible either due to the absence of a required object or when attempting to grow furniture. Accurately identifying impossible goals and generalizing this knowledge is crucial for strong performance in the Little-Zoo environment. In this section, we analyze how MAGELLAN handles impossible goals.
Figure 18(a) illustrates the clustering of different categories of impossible goals. The impossible ”Grasp” goals form a compact cluster, while the ”Grow” goals—categorized as ”Grow plants”, ”Grow herbivores”, ”Grow carnivores”, and ”Grow furniture”—exhibit four less-defined clusters. Additionally, a large, mixed cluster contains various impossible ”Grow” goals. However, when examining the same embeddings through the lens of missing objects, as shown in Figure 18, a distinct structure emerges: MAGELLAN also organizes goals based on the absent object in the scene.
For instance, the previously observed clusters of impossible ”Grow” goals largely align with the ”missing water” category, as water is a prerequisite for all ”Grow” goals. Moreover, the central mixed cluster from Figure 18(a) is further clustered into sub-clusters based on the missing object. In particular, the red cluster at the bottom of Figure 18 represents the absence of a plant, encompassing both ”Grow herbivores” and ”Grow carnivores” goals. This suggests that MAGELLAN effectively captures underlying environmental dependencies to determine goal feasibility. This ability to infer structural properties of the environment likely contributes to the strong generalization performance observed in Section 4.3 and Section 4.4.
Figure 18: Plot of the t-SNE-projected embeddings of impossible goals, where each goal is color-coded based (a) on the type of impossible goal (b) on the specific missing object required to make it possible.
@@
    • High LP (figures 19(b)19(c)19(d)): the agent is getting a high ALP as it is learning some ”Grow carnivores” goals. Here, MAGELLAN outperforms baselines by generalizing its ALP estimation and continuing training on these goals.
    • Medium LP at the end of training (figures 19(e) to  19(j)): the agent has almost learned all the goals, and it only remains few goals in the ”Grow carnivores” category on which its performance is not stable. When changing the goal space, Online-ALP and Uniform take a lot of steps to find the remaining goals with LP in the new goal space, destabilizing the agent. As a result, the agent’s SR decreases just after the transition and does not recover after $50$k training epispdes. MAGELLAN and EK-Online-ALP maintain their high SR.

### v3

@@ D.3.2 Goal sampling strategies
We analyze the sampling strategies of the different methods. In Figure 19, it appears that MAGELLAN has a sampling strategy close to the one of EK-Online-ALP. The main difference is that EK-Online-ALP does not select impossible goals thanks to the use of expert knowledge. MAGELLAN’s sampling strategy quickly begins to sample goals of type ”Grow herbivore” after sampling goals of type ”Grow plant”, which contrasts with the strategy of Online-ALP. Indeed, Online-ALP mostly selects goals already encountered, getting stuck on one goal type before moving to another. The similarity between MAGELLAN and EK-Online-ALP indicates that MAGELLAN is able to cluster the goal space dynamically, using this information to generalize its competence estimation to sample interesting goals not yet discovered. More results on the dynamic clustering of the goal space by MAGELLAN are given in Appendix D.4.2.
@@ D.4.1 Per-goal Success Probability estimation on test set
In this section, we provide a detailed analysis of the results presented in Section 4.3, specifically examining the ability to generalize success probability estimations to test goals across different goal types (see Figure 20). As expected, Online-ALP exhibits the largest errors, as it can only assign a success probability of 0 to unseen goals. In contrast, both MAGELLAN and EK-Online-ALP achieve highly accurate estimations for ”Grasp”, ”Grow plant”, and ”Grow herbivore”. However, both methods tend to overestimate the generalization abilities of the policy on ”Grow carnivore” goals.
@@ D.4.2 Evolution of the embeddings
  • After mastering the goals of type ”Grasp”, Figure 21(b) shows several clustered appeared. All ”Grasp” goals are in the same cluster with a high estimated success probability zone. The goals of type ”Grow plant” are also clustered together and close to the zone of high estimated success probability. That is a hint that MAGELLAN has already picked them as the next candidate type for the curriculum. It also correctly places the ”Grow plant” goals from the test set in the same cluster as the ones from the train set. However, it still mixes them with impossible goals, underlying that it does not fully master this type of goal. The ”Grow herbivore” and ”Grow carnivore” are mixed with other impossible goals.
  • After achieving mastery of ”Grow plant”, in Figure 21(c), we see both ”Grasp” and ”Grow plant” are correctly clustered in the zone of high success probability, with the goals from the test set correctly placed into the two clusters. The goals from the type ”Grow herbivore” and ”Grow carnivore” are clustered together and almost separated from the impossible goals. Their cluster is close to the zone of high estimated probability.
  • Once ”Grow herbivore” has been solved, we see in Figure 21(d) that the three achieved goal types are clustered in three groups, all in the zone of high estimated success probability. The ”Grow carnivore” are far from this zone but still clustered together, apart from the impossible ones. It also appears that the model places ”Grow carnivore” goals from the test close to the ”Grow herbivore”.
  • Finally, when the type ”Grow carnivore” is mastered, in Figure 21(e), all the goals from the possible goal types (from both train and test) are placed in the zone of high success probability. The goals of type ”Grow X” form three clusters tightly packed. The ”Grasp” type is separated from the other types.
@@
(c) ”Grow plant” type of goals mastered.
(d) ”Grow herbivore” type of goals mastered.
(e) ”Grow carnivore” type of goals mastered.
@@ D.4.3 Embedding of impossible goals
As detailed in Section A.3, a goal is considered impossible either due to the absence of a required element or when attempting to grow furniture. Accurately identifying impossible goals and generalizing this knowledge is crucial for strong performance in the Little-Zoo environment. In this section, we analyze how MAGELLAN handles impossible goals.
Figure 22(a) illustrates the clustering of different categories of impossible goals. The impossible ”Grasp” goals form a compact cluster, while the ”Grow” goals—categorized as ”Grow plant”, ”Grow herbivore”, ”Grow carnivore”, and ”Grow furniture”—exhibit four less-defined clusters. Additionally, a large, mixed cluster contains various impossible ”Grow” goals. However, when examining the same embeddings through the lens of missing elements, as shown in Figure 22, a distinct structure emerges: MAGELLAN also organizes goals based on the absent element in the scene.
For instance, the previously observed clusters of impossible ”Grow” goals largely align with the ”missing water” category, as water is a prerequisite for all ”Grow” goals. Moreover, the central mixed cluster from Figure 22(a) is further clustered into sub-clusters based on the missing element. In particular, the red cluster at the bottom of Figure 22 represents the absence of a plant, encompassing both ”Grow herbivore” and ”Grow carnivore” goals. This suggests that MAGELLAN effectively captures underlying environmental dependencies to determine goal feasibility. This ability to infer structural properties of the environment likely contributes to the strong generalization performance observed in Section 4.3 and Section 4.4.
Figure 22: Plot of the t-SNE-projected embeddings of impossible goals, where each goal is color-coded based (a) on the type of impossible goal (b) on the specific missing element required to make it possible.
@@
  • Scenario zero LP (Figure 23(a)): the agent has mastered the ”Grasp” and ”Grow plant” goals and has 00 ALP across all goals. In this scenario, all ALP estimations are equivalent. EK-Online-LP manages to discover new ALP niches faster as all impossible goals are in the same group.
@@
    • High LP (figures 23(b)23(c)23(d)): the agent is getting a high ALP as it is learning some ”Grow carnivore” goals. Here, MAGELLAN outperforms baselines by generalizing its ALP estimation and continuing training on these goals.
    • Medium LP at the end of training (figures 23(e) to  23(j)): the agent has almost learned all the goals, and it only remains few goals in the ”Grow carnivore” category on which its performance is not stable. When changing the goal space, Online-ALP and Uniform take a lot of steps to find the remaining goals with LP in the new goal space, destabilizing the agent. As a result, the agent’s SR decreases just after the transition and does not recover after $50$k training epispdes. MAGELLAN and EK-Online-ALP maintain their high SR.